# Hierarchical Open-Vocabulary 3D Scene Graphs for Language-Grounded Robot Navigation

### v1

@@ III-C Object Navigation from Long Queries
HOV-SG extends the scope of potential navigation goals to more specific spatial concepts like regions and floors compared to simple object goals [11, 16, 7, 10]. Language-guided navigation with HOV-SG involves processing complex queries such as “find the toilet in the bathroom on floor 2” using a large language model (GPT-3.5). We break down such lengthy instructions into three separate queries: one for the floor level, one for the room level, and one for the object level. Leveraging the explicit hierarchical structure of HOV-SG, we sequentially query against each hierarchy to progressively narrow down the solution space. This is done by taking the cosine similarity between the identified query floor, region and object as well as all objects, rooms and floors given in the graph, respetively. Once a target node is identified via scoring, we utilize the navigational graph mentioned above to plan a path from the starting pose to the target destination, which is demonstrated in Fig. S.1 and visualized in Fig. 1.
@@ IV Experimental Evaluation
In the following, we first evaluate HOV-SG against recent open-vocabulary map representations on the task of 3D semantic segmentation on the ScanNet [41] and Replica datasets [42] in Sec. IV-A. In the same section, we justify the chosen approach by ablating several of our key contributions on the Replica dataset. Secondly, we investigate the accuracy of the constructed hierarchical, open-vocabulary 3D scene graphs on scenes of the Habitat Matterport 3D Semantics Dataset [43] (HM3DSem) in Sec. IV-B. Third, we study the open-vocabulary object retrieval capabilities of HOV-SG and demonstrate large-scale language-grounded robotic navigation in the real world (Sec. IV-C). In addition, we also compare the memory footprint of sparse and dense map representations in Sec. IV-D.
@@
We evaluate the open-vocabulary 3D semantic segmentation performance of HOV-SG on the ScanNet [41] and Replica [42] datasets. We compare our method with two competitive vision-and-language representations, namely ConceptFusion [10] and ConceptGraphs [11], and ablate over two CLIP backbones, see Table I. We consider the ViT-H-14 and a fine-tuned ViT-L-14 CLIP backbone released with the work OVSeg [39]. For each category in the dataset, we use the template: “There is the category in the scene.“ and then generate its CLIP text embeddings [20] and the category name itself. Next, we average them to obtain the embedding of the specific category. We obtain predicted labels for each object node by computing the cosine similarity between all object nodes’ embeddings and all category embeddings and lastly apply the operator.
@@
Then we accumulate all objects’ point clouds to create our predicted point cloud and transform the point cloud to the same coordinate frame as that of the point cloud with ground-truth (GT) semantic labels . Given that the predicted point cloud may exhibit varying densities compared to the ground truth (GT), we iterate through each GT point to locate its five nearest points in , and then determine the majority label among these points as the predicted label wrt. that respective GT point.
@@
For consistency, we evaluate the same scenes evaluated in [11, 10]. For ScanNet, we evaluate scenes: scene0011_00, scene0050_00, scene0231_00, scene0378_00, scene0518_00. For Replica, we evaluate the following scenes: office0-office4 and room0-room2.
@@
The semantic segmentation results on both Replica and ScanNet are shown in Tab. I. In terms of mIOU and F-mIOU, HOV-SG outperforms the open-vocabulary baselines by a large margin. This is primarily due to the following improvements we made: First, when we merge segment features, we consider all point-wise features that each segment covers and use DBSCAN to obtain the dominant feature, which increases the robustness compared to taking the mean as done by ConceptGraphs. Secondly, when we generate the point-wise features, we use the mask feature which is the weighted sum of the sub-image and its contextless counterpart. This mitigates the impact of salient background objects. Further qualitative results are shown in Fig. 6.

### v2

@@ III-C Navigation with Scene Graph
HOV-SG extends the scope of potential navigation goals to more specific spatial concepts such as regions and floors compared to simple object goals [14, 16, 10, 13]. Language-guided navigation with HOV-SG involves processing complex queries such as “find the toilet in the bathroom on floor 2” using a large language model (prompts are given in the supplementary material Sec. S.1-A). We break down such lengthy instructions into three separate queries: one for the floor level, the room level, and the object level, respectively. Leveraging the explicit hierarchical structure of HOV-SG, we sequentially query against each hierarchy level to progressively narrow down the solution corridor. This is done by taking the cosine similarity between the identified query floor, query region, and query object as well as all objects, rooms, and floors given in the graph, respectively. Once a target node is identified via scoring, we utilize the navigational graph mentioned above to plan a path from the starting pose to the target destination, which is demonstrated in Fig. S.1 and visualized in Fig. 1.
@@ IV Experimental Evaluation
The goals of our experiments are five-fold: (i) we quantitatively compare HOV-SG with recent open-vocabulary map representations in 3D semantic segmentation on ScanNet and Replica (Sec. IV-A), (ii) we investigate the semantic and geometric accuracy of HOV-SG at the floor, room, and object level on the Habitat Matterport 3D Semantic Dataset (Sec. IV-B), (iii) we study how HOV-SG enables large-scale language-grounded navigation in the real-world (Sec. IV-C), (iv) we demonstrate the compact memory footprint of HOV-SG compared to previous open-vocabulary representations (Sec. IV-D), and lastly, (v) we justify our design choices through an ablation study (Sec. IV-E).
@@
Higher values are better. The used evaluation metrics are defined in Sec. S.2-B in the supplementary material. The ConceptFusion pipeline evaluated against made use of instance masks predicted by SAM [22]. The MinkowskiNet [50] is a privileged method that was trained on the full set of ScanNet [24] scenes to demonstrate the gap between zero-shot and fully-supervised methods.
@@
To test the semantic expressiveness of our HOV-SG method, we evaluate the open-vocabulary 3D semantic segmentation performance on ScanNet [24] and Replica [23]. We compare our method with two alternative vision-language representations (ConceptFusion [13] and ConceptGraphs [14]) while using different CLIP backbones. We consider ViT-H-14 and a fine-tuned backbone ViT-L-14 released with the work OVSeg [48]. The used evaluation metrics are defined in Sec. S.2-B. To demonstrate the existing gap between zero-shot and fully supervised methods we also evaluated a MinkowskiNet [50] instance trained on ScanNet [24].
@@
Prediction Generation: We generate the CLIP text embedding for each category contained in the dataset by using a template of the form “There is the {category} in the scene.” as well as the category name “{category}” itself. Next, we average the two to obtain the embedding of each specific category. We obtain predicted labels for each object node by computing the cosine similarity between all object nodes’ embeddings and all category embeddings and lastly apply the operator. In the following, we concatenate all objects’ point clouds to create our predicted point cloud and transform it to the same coordinate frame as that of the point cloud with ground-truth (GT) semantic labels . Given that the predicted point cloud may exhibit varying point densities compared to the ground truth (GT), we iterate through each GT point to locate its five nearest points in , and then determine the majority label among these points as the predicted label for each GT point.
@@
Evaluation Scenes: For consistency, we evaluate the same scenes evaluated in [14, 13]. For ScanNet, we evaluate scenes: scene0011_00, scene0050_00, scene0231_00, scene0378_00, scene0518_00. Regarding Replica, we evaluate on office0-office4 and room0-room2.
@@
Results: The semantic segmentation results on both Replica and ScanNet are provided in Table I. Regarding mIOU and F-mIOU, HOV-SG outperforms the open-vocabulary baselines by a large margin. This is primarily due to the following improvements we made: First, when we merge segment features, we consider all point-wise features that each segment covers and use DBSCAN to obtain the dominant feature, which increases the robustness compared to taking the mean as done by ConceptGraphs [14]. Secondly, when we generate the point-wise features, we use the mask feature which is the weighted sum of the sub-image and its contextless counterpart. This mitigates the impact of salient background objects. Further qualitative results are given in Fig. 6.
IV-B Scene Graph Evaluation on Habitat 3D Semantics
We evaluate our scene graph on four aspects. To analyze the geometric accuracy of the scene graph, we analyze the floor and class-agnostic region segmentation performance in Sec. IV-B1. To evaluate the semantic accuracy, we evaluate predicted region semantics (Sec. IV-B2) as well as open-vocabulary object-level semantics (Sec. IV-B3). To scrutinize the downstream navigation capabilities of HOV-SG, we conduct hierarchical object retrieval and navigation experiments given abstract language queries in Sec. IV-B4. We display two exemplary constructed hierarchical 3D scene graphs in Fig. 7. The 3D scene graph visualization of the remaining scenes is shown in Fig S.4.
Dataset: In order to evaluate various aspects of the produced scene graph hierarchy, we have chosen the Habitat-Semantics dataset (HM3DSem) as it provides true open-vocabulary labels across large multi-floor scenes and also provides object-region assignments. Since our approach operates on RGB-D frames, we manually record random walks of 8 scenes of the Habitat Semantics dataset [25], which span multiple rooms and floors: 00824, 00829, 00843, 00861, 00862, 00873, 00877, 00890. To construct ground-truth maps to compare against, we fuse the RGB-D and panoptic data across all frames given accurate odometry and obtain RGB and panoptic global ground truth. These maps are finally voxelized using a resolution.
IV-B1 Floor and Class-Agnostic Room Segmentation
In order to evaluate both floor and class-agnostic room segmentation, we identified several heuristics on top of the provided metadata of the dataset. Since HM3DSem does not include floor height labels, we hand-labeled all upper and lower floor boundaries. In addition, we pooled the annotated objects contained in our constructed ground truth point clouds based on their associated region labels. Based on this, we obtain region-wise point clouds we utilize as ground truth. As shown in Table II, our method achieves 100% accuracy in retrieving the number of floors, both in single-floor as well as multi-floor scenes. In addition, we evaluate the region segmentation performance of HOV-SG and compare with Hydra [19] across eight scenes on HM3DSem. We observe slightly lower precision but a significantly greater recall of 83.59% compared to 77.55%. In addition to the overall results given in Table II, we provide scene-wise results in Table S.1 in the supplementary material. In general, we obtain higher precision and recall on smaller scenes comprising fewer regions. Similar to Hydra [19], our approach utilizes a naïve morphological heuristic to segregate regions, which does not work well on more complex, semantically ambiguous room layouts such as combined kitchen and living rooms. Nonetheless, our approach does not suffer from this drawback too drastically as we equip each segmented region with 10 representative embeddings. This allows adaptive prompting without directly setting a fixed room category.
| Method | Floors | Regions | |
| [%] | Precision [%] | Recall [%] | |
| Hydra [19] | - | 86.18 | 77.55 |
| HOV-SG (ours) | 100 | 84.10 | 83.59 |
Evaluation of the floor and room segmentation: We present the accuracy of correctly predicted floors using a threshold of . The region segmentation precision (P) and recall (R) are calculated based on the metric in Hydra [19].
| (a) Scene 00824 (single floor) | (b) Scene 00861 (two floors) |
IV-B2 Semantic Room Classification
We quantitatively evaluate our proposed view embedding-based room category labeling method (See Fig. 4) by comparing it against two strong baselines across the set of eight scenes on HM3DSem. Both baselines rely on object labels to classify room categories. To draw a fair comparison, all methods rely on ground truth room segmentation, namely the class-agnostic mask of each room. Thus, all objects are assigned to rooms based on ground truth room layouts. This is different from the general HOV-SG method, which also estimates room masks.
Dataset: In this evaluation, we utilize a closed set of room categories. The HM3DSem dataset does not provide annotated room categories but merely educated votes, which are often not sufficient. Therefore, we manually labeled the regions of the eight scenes detailed in Sec. IV-B. The list of room categories is provided in Sec. S.2-D.
Baselines: We compare the HOV-SG approach of using filtered view embeddings for labeling rooms against a privileged and an unprivileged baseline. The privileged baseline operates on ground truth object categories contained within each room. In order to obtain room labels, the baseline prompts an LLM (GPT-3.5 and GPT-4 [43]) to provide a room category guess out of the closed set of room categories given the objects per room in a few-shot manner (prompts are detailed in Sec. S.2-D). The second and unprivileged baseline applies the same principle of prompting an LLM but only has access to the predicted object categories obtained using HOV-SG. This means that each predicted object is labeled as the category showing the highest similarity to the object feature among a category list. In general, we expect that the number of objects will be different from the privileged baseline because of under- and over-segmentation of HOV-SG’s predictions. In comparison, our view embedding method relies on 10 distinct view embeddings, which are scored against the chosen set of room categories. The final predicted room category is defined by the room category that showed the highest similarity across all view embeddings as described in Fig. 4.
Metrics: We apply two different evaluation criteria: The first accuracy called fosters replicability by evaluating whether the predicted and the ground truth room categories are text-wise equal. Different from that, the performance regarding the metric is produced via human evaluation. This is crucial as room categories are not always fully determinable when labeling, e.g., combined kitchen and living room areas. Moreover, the answers provided by the LLM do not always state definitive categories because of frequent hallucinations. A high number of objects per room exacerbates this. This particularly applies to the unprivileged baselines when facing under-segmentation. In order to circumvent this, we manually filter all outputs across the set of eight scenes and check whether the LLM leaned towards the correct answer such as predicting a synonym of the GT room type, which boosts results in favor of the LLM-based methods. In addition to this, we also evaluated the same task using the current state-of-the-art LLM, GPT-4, which shows significantly fewer hallucinations and increased accuracy.
Results: As presented in Table III, the view embedding method of HOV-SG outperforms all unprivileged baselines both in terms of the strict accuracy () as well as the approximate accuracy () by a significant margin. In addition, HOV-SG also outperforms the privileged baseline relying on GPT-3.5 while showing similar approximate accuracy as GPT-4 of 84.10% vs. 84.25%. Thus, we conclude that our room labeling method is robust and outperforms comparable unprivileged methods by a significant margin. Furthermore, we provide additional scene-wise evaluations in Table S.2 in the supplementary material.
| Room Identification Method | [%] | [%] | |
| Privileged | GPT-3.5 w/ GT | 66.89 | 81.49 |
| object categories | |||
| GPT-4 w/ GT | 79.86 | 84.25 | |
| object categories | |||
| Unprivileged | GPT-3.5 w/ predicted | 28.48 | 42.95 |
| object categories | |||
| GPT-4 w/ predicted | 59.47 | 62.55 | |
| object categories | |||
| HOV-SG (ours) | 73.93 | 84.10 | |
| w/ view embeddings | |||
We present the room classification performance on HM3DSem of HOV-SG (view embeddings) and two baselines (GPT-3.5 / GPT-4) using either privileged or unprivileged information. In the privileged case, rooms are labeled based on ground truth object categories per room. The unprivileged baselines take the predicted masks and categories as input. We consider two different evaluation criteria: measures whether the exact text-wise room category was predicted while measures semantically correct room predictions based on qualitative human evaluation through manual inspection.
IV-B3 Object-Level Semantics
Existing open-vocabulary evaluations usually circumvent the problem of measuring true open-vocabulary semantic accuracy. This is due to arbitrary sizes of the investigated label sets, a potentially enormous and challenging amount of object categories [25], and the ease of use of existing evaluation protocols [10, 13]. While human-level evaluations such as Amazon Mechanical Turk (AMT) partly solve this problem, robust results replication and scaling remain challenging [14].
Metrics: In this work, we propose the novel AUC metric that quantifies the area under the top- accuracy curve between the predicted and the actual ground-truth object category (see Fig. S.2). This entails computing the ranking of all cosine similarities between the predicted object feature and all possible category text features, which are in turn encoded using a vision-language model (CLIP). Thus, the metric encodes how many erroneous shots are necessary on average before the ground-truth label is predicted correctly. Based on this, the metric encodes the actual open-set similarity while scaling to large, variably-sized label sets. We envision a future use of this metric in diverse open-vocabulary tasks.
We visualize the AUC curve HOV-SG on the 00824 scene of HM3DSem in Fig. S.2. The closer the curve is to the upper left corner of the plot, the higher the open-vocabulary similarity. Instead of showing the accuracy at distinct values of , we normalize over the extent of the label set, which contains 1624 categories for HM3DSem. This also shows visually how the AUC metric provides a dependable measure for large but variably sized label sets.
Baselines: In order to show the applicability of the AUC metric, we compare HOV-SG against two strong baselines VLMaps [10] and ConceptGraphs [14] on the Habitat-Semantics dataset [25], which comprises an enormous label set of 1624 object categories. To allow for a fair comparison, we perform a linear assignment among predicted and GT objects and only consider predicted objects that show an with the ground truth. Since VLMaps [10] does not predict masks by design, it takes the object masks predicted by HOV-SG and averages all masked voxels’ features as the object feature.
Results: The overall AUC score as well as various top-k thresholds are given in Table IV. We observe that VLMaps [10] performs inferior, which is presumably due to its dense feature aggregation as well as its dependence on a fine-tuned VLM, LSeg [51], limiting its generalization in challenging open-vocabulary scenarios. It does not only score merely 5% of objects correctly given its top-5 choices but also only predicts the correct class given its top-500 predicted classes in 40.02% of all cases. In comparison, ConceptGraphs [14] obtains a competitive score of 84.07% while HOV-SG achieves 84.88%. Especially, up to the top-100 highest ranking classes, HOV-SG outperforms ConceptGraphs.
| Method | AUC | ||||||
| VLMaps [10] | 0.05 | 0.17 | 0.54 | 15.32 | 26.01 | 40.02 | 56.20 |
| ConceptGraphs [14] | 18.11 | 24.01 | 33.00 | 55.17 | 70.85 | 81.55 | 84.07 |
| HOV-SG (ours) | 18.43 | 25.73 | 36.41 | 56.46 | 69.95 | 80.86 | 84.88 |
We provide object-level semantic accuracies across all eight considered scenes within HM3DSem using both the overall AUC metric across 1624 categories as well as accuracies at a few selected thresholds . To allow for a fair comparison, we perform a linear assignment among predicted and GT objects and only consider predicted objects that show an with the ground truth. Since VLMaps [10] does not predict masks by design, it takes the masks predicted by HOV-SG and evaluates wrt. those.
IV-B4 Hierarchical Concept Retrieval
To take advantage of the hierarchical character of our proposed representation, we evaluate to what extent we can retrieve objects from hierarchical queries of the form: “pillow in the living room on the second floor” or “bottle in the kitchen”. To do so, we decompose the query using GPT-3.5 into its sought-after hierarchical concepts, e.g., [floor 2, living room, pillow] or [-, kitchen, bottle]), and compute the corresponding CLIP embeddings. In the next step, we hierarchically query against the most suitable floor, the most appropriate room, and lastly, the most suitable object given the query at hand (see Table V). While floor prompting is done naively, we select the room producing the highest maximum cosine similarity to the query room across its ten embeddings. On average, this produces higher success rates compared with mean- or median-based schemes.
Retrieval Results: In the following, we compare HOV-SG against an augmented variant of ConceptGraphs [14] that is equipped with privileged floor information and it scores objects against the requested room and object, which allows it to draw answers at the floor and room level. As shown in Table V, HOV-SG shows a significant performance increase of 11.69% on object-room-floor queries and a 2.2% advantage on object-room queries when compared with ConceptGraphs. While ConceptGraphs struggles on larger scenes and under more detailed queries, HOV-SG outperforms it by a significant margin even though it suffers from erroneous room segmentations by design. For more information, we refer to Sec. S.1-A.
| Query Type | Method | # Trials | Retrieval-SR10 [%] | Navigation-SR [%] |
| (o, r, f) | ConceptGraphs | 40.63 | 16.31 | - |
| HOV-SG (ours) | 28.00 | 37.32 | ||
| (o, r) | ConceptGraphs | 34.88 | 29.26 | - |
| HOV-SG (ours) | 31.48 | 40.41 |
Evaluation over 20 frequent distinct object categories in terms of the top-5 accuracy. A match is counted as a success when the between predicted object and ground truth. The number of trials is an average number of trials across the eight scenes evaluated. It is lower for (o,r) compared to (o,r,f) due to a higher number of query duplicates whenever we drop the floor specification. The 20 evaluated categories are: picture, pillow, door, lamp, cabinet, book, chair, table, towel, plant, sink, stairs, bed, toilet, tv, desk, couch, flowerpot, nightstand, faucet.
Language-Grounded Navigation in Simulation: In addition, to the general querying tasks we also perform navigational trials using the Habitat Simulator. Based on the obtained actionable multi-floor navigational Voronoi graphs introduced in Sec. III-B, we traverse the set of retrieved high-probability objects satisfying the query and report the physical retrieval success rate. If the robot stops by one of the matched ground truth point clouds and its distance to it is smaller than , we consider the navigation successful. As shown in Table V, the navigational success rates are higher when compared to the retrieval success rates. We found that the robot regularly reaches the locations of the predicted objects that are in close vicinity to actual ground truth target objects. This effect is measurable since we rely on a Euclidean distance-based evaluation for the navigational success assessment. Moreover, imperfectly predicted instance masks increase the chance of retrieving segments that do not provide complete overlap with actual ground truth objects in terms of geometry and semantics, which induces slight offsets in the retrieved positions when querying. Example navigation trials are shown in Fig. 8.
| Query | # | Graph Querying | Goal Navigation | ||
| Type | Trials | # Successes | SR [%] | Success | SR [%] |
| Object | 41 | 29 | 70.7 | 23 | 56.1 |
| Room | 9 | 5 | 55.6 | 5 | 55.6 |
| Floor | 2 | 2 | 100 | 2 | 100 |
IV-C Real World Language-Grounded Navigation
To validate the system in the real world, we utilize a Boston Dynamics Spot robot with a calibrated Azure Kinect RGB-D camera and a 3D LiDAR attached to it. We collect a stream of RGB-D sequences inside a two-story office building, traversing through a variety of rooms with diverse semantic information as shown in Fig. 9.
Real-World Application of HOV-SG: We calibrate the extrinsics between the LiDAR and the RGB-D camera and apply an off-the-shelf LiDAR SLAM implementation that combines FAST-LIO2 [52] with a loop closure component to obtain LiDAR poses. Subsequently, we leverage the extrinsics to derive the associated camera poses. Finally, employing the RGB-D data and odometry, we construct the HOV-SG representation as detailed in Sec. III.
Robot Navigation with Long Queries: Within the two-story building, we select 41 object goals, nine room goals, and two floor goals and use natural language to query the HOV-SG representation built. Some examples of the queries are “go to floor 0”, “navigate to the kitchen on floor 1”, and “find the plant in the office on floor 0”. Unlike previous methods that only retrieved object-level goals, our representation enables long queries containing multiple levels of concepts and facilitates a more detailed constrained retrieval.
To separate the evaluation of our representation and the navigation system, we first evaluate the accuracy of the retrieval tasks qualitatively. Since the building is well structured, we can easily determine the boundary between rooms and regions like offices, corridors, and dining rooms. Meanwhile, we manually label the categories of all regions in the building, using a category set containing office, corridor, kitchen, seminar room, meeting room, dining room, bathroom. If the HOV-SG representation returns the correct floor and room point cloud as well as an object point cloud that shows overlap with the correct object, we consider this as retrieval success. We achieve a 100% success rate in floor retrieval, a 55.6% success rate in room retrieval, and a 70.7% success rate in object retrieval. The major failure cases for room retrieval stem from the visual similarity among rooms such as “meeting room”, “seminar room”, and “dining room” as shown in Fig. S.3.
We query our HOV-SG representation using hierarchical concepts and utilize the Spot quadrupedal robot to carry out navigation trials. In our experiments, the robot navigates to queried objects across both floors with a 56.1% success rate while navigating to all successfully retrieved room and floor concepts with language instructions. One failure case occurred for the “go to the whiteboard in the office on the second floor” query. Since the whiteboard is attached to a room-separating wall, the robot achieved the necessary distance to the object but was positioned on the opposite side of the wall as shown in Fig. S.6). In addition to that, we did not consider target locations on stair segments to prevent the robot assuming unstable poses. The evaluation results for retrieval and navigation are shown in Table VI. We display a subset of target objects in Fig. S.6 and three trials in more detail in the supplementary material Fig. S.5.
IV-D Representation Storage Overhead Evaluation
A key advantage of HOV-SG is the compactness of the representation. We compare the storage size of VLMaps [10], ConceptGraphs [14], and HOV-SG created for the eight scenes in the HM3DSem dataset and show the results in Tab. VII. We adapt VLMaps to store LSeg features at 3D voxel locations with 0.05m voxel size. The backbone of the LSeg is ViT-B-32, which has 512-dimensional features. ConceptGraphs and HOV-SG are using the ViT-H-14 CLIP backbones, which requires saving 1024-dimension features in the representation. VLMaps is optimized to only save features at voxels near object surfaces instead of saving redundant features at non-occupied voxels. Nonetheless, thanks to the compact graph structure, ConceptGraphs and HOV-SG are much smaller than their dense counterparts. HOV-SG even reduces as much as 75% in memory footprint on average compared to VLMaps. While ConceptGraphs encodes supplementary object relationships, HOV-SG incorporates hierarchical semantic features. Overall, ConceptGraphs and HOV-SG serve as excellent complementary representations, each emphasizing distinct facets of scene semantics.
IV-E Ablation Study
In order to shed light on the contributions of various key components in our approach, we present an ablation study on the Replica dataset [23] in Table VIII. One key component of the 3D open-vocabulary segment-level mapping pipeline (Sec. III-A) is the DBSCAN clustering we apply to pixel-wise CLIP embeddings associated with each segment to select the most representative features among them. This design, inspired by the principle of majority voting, has proven effective in mitigating outlier CLIP features caused by the inherent limitations of CLIP and the noise originating from SAM’s outputs, thereby enhancing semantic accuracy. A different key component of our approach involves fusing CLIP features extracted from various sources: the global image, the masked image of an object based on its SAM mask, and the masked object image given the SAM mask without background. In contrast to ConceptGraphs [14], which only integrate the global image embedding and the sub-image embedding, we hypothesize that incorporating salient features from the sub-image into CLIP embeddings could enhance accuracy. Based on this, we tested three setups: utilizing only the CLIP embedding of the masked object including background (L-CLIP), employing only the CLIP embedding of the masked object without background (M-CLIP), and third, combining both of these CLIP embeddings by fusing them as done in HOV-SG. Our findings indicate that combining both embeddings yields the highest semantic accuracy as given in Table VIII.